# thanks: This work is based on the author’s M.Sc. dissertation (Ref.[1]), but further correcting and expanding upon it

### v1

@@ VI.3 Weyl Semimetals and Beyond
The electroaxial effective action derived above has direct applications to time-reversal asymmetric Weyl and Dirac semimetals, where axial-vector backgrounds emerge from Weyl-node separation $b_{\mu}=(b_{0},\mathbf{b})$ in momentum space, strain-induced pseudomagnetic fields, and magnetic textures [31, *Diractorsion2, *Diractorsion, 96, *torsionweylmetal, *EMWeylsemimetals]. In this context, the node separation vector plays the role of $S_{\mu}$, and our nonperturbative results become experimentally accessible: the required magnetic field strengths ($\sim 1$1010 T) are readily achievable in laboratory settings. Lorentz invariance is emergent rather than fundamental in these systems [35], permitting nonzero mixed terms $S_{\mu\nu}F^{\mu\nu}=\mathbf{E}_{A}\cdot\mathbf{B}_{S}+\mathbf{E}_{S}\cdot\mathbf{B}_{A}\neq 0$ that induce apparent vector-current nonconservation, requiring Bardeen-Zumino polynomial counterterms for consistent anomaly structure.
Our framework yields several testable predictions. First, the complete chiral-current divergence [Eqs. (82)]—which includes contributions beyond the standard Adler–Bell–Jackiw / Carroll–Field–Jackiw-Torsion [Eqs. (73),(74)] and Nieh–Yan terms—modifies anomalous transport coefficients and should manifest as corrections to longitudinal magnetoconductance in materials such as TaAs or Cd3As2. Second, the pair-production rates derived in Eqs. (91)–(94) translate into chiral-charge pumping rates between Weyl nodes under strong applied fields, with the vacuum-stabilization effect (Fig. 2) predicting enhanced threshold fields compared to naive Schwinger estimates, this effect should be observable in pulsed magnetic field experiments. Third, the electromagnetic-field generation via axial-vector decay [Eq. (75)] suggests that dynamical strain or magnetic textures in Weyl semimetals can source photon production, analogous to phenomena observed in magnetized plasmas [98, *plasma_axial2]. Quantitative comparison with transport measurements and pump-probe spectroscopy will be pursued in subsequent work.
@@ VII Discussion and Outlook
Finally, our treatment is confined to flat spacetime. Gravitational effects, which are crucial for cosmological applications like baryogenesis or primordial magnetogenesis, are thus incorporated only indirectly via the axial field’s origin (e.g., as torsion’s pseudo-trace). A fully covariant generalization, incorporating curved backgrounds [107, *HeatKernelQG2] and spin-connection couplings, would be essential to assess backreaction, gravitational particle production [109], and the interplay between spacetime curvature and axial-induced anomalies [110].
@@ Appendix A Detailed Expressions of the Auxiliary Terms
$\Delta=2\mathcal{K}+i\sqrt{2\mathcal{K}+64\left\{4\left(\mathcal{A}^{2}-\eta^{2}\Pi_{\mu}S^{\mu}\mathcal{V}\right)+3\mathcal{W}^{2}+12\eta^{2}m^{2}\left[\eta^{2}\left(\Pi_{\mu}S^{\mu}\right)S^{2}-\left(\tilde{V}^{\mu\nu}S_{\nu}\right)^{2}\right]\right\}^{3}}$ (104)

### v2

@@ VI.3 Weyl Semimetals and Beyond
The electroaxial effective action derived above has direct applications to time-reversal asymmetric Weyl and Dirac semimetals, where axial-vector backgrounds emerge from Weyl-node separation $b_{\mu}=(b_{0},\mathbf{b})$ in momentum space, strain-induced pseudomagnetic fields, and magnetic textures [40, *Diractorsion2, *Diractorsion, 103, *torsionweylmetal, *EMWeylsemimetals]. In this context, the node separation vector plays the role of $S_{\mu}$, and our nonperturbative results become experimentally accessible: the required magnetic field strengths ($\sim 1$1010 T) are readily achievable in laboratory settings. Lorentz invariance is emergent rather than fundamental in these systems [44], permitting nonzero mixed terms $S_{\mu\nu}F^{\mu\nu}=\mathbf{E}_{A}\cdot\mathbf{B}_{S}+\mathbf{E}_{S}\cdot\mathbf{B}_{A}\neq 0$ that induce apparent vector-current nonconservation, requiring Bardeen-Zumino polynomial counterterms for consistent anomaly structure.
Our framework yields several testable predictions. First, the complete chiral-current divergence [Eqs. (85)]—which includes contributions beyond the standard Adler–Bell–Jackiw / Carroll–Field–Jackiw-Torsion [Eqs. (76),(77)] and Nieh–Yan terms—modifies anomalous transport coefficients and should manifest as corrections to longitudinal magnetoconductance in materials such as TaAs or Cd3As2. Second, the pair-production rates derived in Eqs. (94)–(97) translate into chiral-charge pumping rates between Weyl nodes under strong applied fields, with the vacuum-stabilization effect (Fig. 2) predicting enhanced threshold fields compared to naive Schwinger estimates, this effect should be observable in pulsed magnetic field experiments. Third, the electromagnetic-field generation via axial-vector decay [Eq. (78)] suggests that dynamical strain or magnetic textures in Weyl semimetals can source photon production, analogous to phenomena observed in magnetized plasmas [105, *plasma_axial2]. Quantitative comparison with transport measurements and pump-probe spectroscopy will be pursued in subsequent work.
@@ VII Discussion and Outlook
Finally, our treatment is confined to flat spacetime. Gravitational effects, which are crucial for cosmological applications like baryogenesis or primordial magnetogenesis, are thus incorporated only indirectly via the axial field’s origin (e.g., as torsion’s pseudo-trace). A fully covariant generalization, incorporating curved backgrounds [114, *HeatKernelQG2] and spin-connection couplings, would be essential to assess backreaction, gravitational particle production [116], and the interplay between spacetime curvature and axial-induced anomalies [117].
@@ Appendix A Detailed Expressions of the Auxiliary Terms
$\Delta=2\mathcal{K}+i\sqrt{2\mathcal{K}+64\left\{4\left(\mathcal{A}^{2}-\eta^{2}\Pi_{\mu}S^{\mu}\mathcal{V}\right)+3\mathcal{W}^{2}+12\eta^{2}m^{2}\left[\eta^{2}\left(\Pi_{\mu}S^{\mu}\right)S^{2}-\left(\widetilde{V}^{\mu\nu}S_{\nu}\right)^{2}\right]\right\}^{3}}$ (107)